OpenStudy (anonymous):

The equations of motion of a point particle in space are given by: (i) x(t) = 2t2 − 3 ; y(t) = −3t2 ; z(t) = 4t2 − 5 (ii) x(t) = asin(ωt); y(t) = acos(ωt); z(t) = bt In each case, calculate the velocity vector and the acceleration vector for an arbitrary value of t (considering a and ω as fixed parameters). Calculate the absolute values of those vectors.

OpenStudy (anonymous):

Motion -> Velocity (Take one derivative of each equation) Then to get to acceleration, take another derivative. (i): \[dx/dt = v(t) = 4t\] \[dy/dt = v(t) = -6t\] \[dz/dt = v(t) = 8t\] \[d^{2}x/dt^{2} = dv/dt = 4\] \[d^{2}y/dt^{2} = dv/dt = -6\] \[d^{2}z/dt^{2} = dv/dt = 8\] (ii): \[dx/dt = \omega*a*\cos(\omega*t)\] \[dy/dt = -\omega*a*sin(\omega*t)\] \[dz/dt = b\] \[d^{2}x/dt^{2} = -\omega^{2}*a*\sin(\omega*t)\] \[d^{2}y/dt^{2} = -\omega^{2}*a*\cos(\omega*t)\] \[d^{2}z/dt^{2} = 0\] For absolute values of these, just drop the (-) sign

OpenStudy (anonymous):

Thank you!